OpenStudy (anonymous):

Factor the following trinomial. w18 – 9w9y5 + 14y10

OpenStudy (anonymous):

Take whatever the things you can take as common..

OpenStudy (anonymous):

Hint: What squared equals w^18?

OpenStudy (anonymous):

No, here you can do one thing more: put\[w^9 = x\] So, the equation becomes: \[x^2 - (9y^5)x + 14y^{10} = x^2 - (7y^5)x - (2y^5)x + 14y^{10}\] \[x(x-7y^5) - 2y^5(x - 7y^5) = (x - 2y^5)(x-7y^5)\]

OpenStudy (anonymous):

Now you know what is substituted as x?? replace x by that value and get the answer...

OpenStudy (anonymous):

Look my high school was voted worst in the tri-state area and half my class failed math so I have no freaking clue what I'm trying to do and I need all the help I can get.

OpenStudy (anonymous):

Thank you waterineyes for helping!

OpenStudy (anonymous):

Welcome dear...